Question no. 01: A. Solve the following Equations: a) |π‘₯ βˆ’ 6| = 3 b) |π‘₯ βˆ’ 4| = |βˆ’3π‘₯ + 8| c) |2π‘₯ + 4| = |5π‘₯ + 2| B. Find the midpoint and distance of the line segment that connects the following points a) (0,3) and (4,7) b) (-9,6) and (-1, -2) c) (-4.5, -12.5) and (4.5, 13) Question no. 02: A. Find the slope, x-intercept and y-intercept form of the following equations a) 5x + 2y =-10 b) 13y -2x = 3 c) 25𝑦 + 31π‘₯ βˆ’ 18 = 10𝑦 d) βˆ’3π‘₯ + 4𝑦 βˆ’ 10 = 7π‘₯ βˆ’ 2𝑦 + 50 B. Find the slope of the line that connects the following points a) (0,3) and (4,7) b) (-9,6) and (-1, -2) C. Determine the slope-intercept form of a linear equation, given the listed attributes: a) Slope = -2 and y-intercept = (0,10) b) Slope = -3 and (4, -2) lies on line c) Slope = 0 and (2,4) lies on line d) (3, -2) and (-12,1) lies on line e) (20, 240) and (15,450) lies on line


1
Expert's answer
2021-07-28T17:51:03-0400

1.

A.

a.

"|x-6|=3"

"x-6=-3\ or\ x-6=3"

"x=3, 9"

b.


"|x-4|=|-3x+8|"

"x-4=-3x+8\ or\ x-4=3x-8"

"4x=12\ or\ 2x=4"

"x=2, 3"



c.


"|2x+4|=|5x+2|"

"2x+4=5x+2\ or\ 2x+4=-5x-2"

"3x=2\ or\ 7x=-6"

"x=-\dfrac{6}{7}, \dfrac{2}{3}"

B.

a.


"x=\dfrac{0+4}{2}=2, y=\dfrac{3+7}{2}=5"

"M(2, 5)"


"l=\sqrt{(4-0)^2+(7-3)^2}=4\sqrt{2}"

b.


"x=\dfrac{-9-1}{2}=-5, y=\dfrac{6-2}{2}=2"

"M(-5, 2)"




"l=\sqrt{(-1-(-9))^2+(-2-6)^2}=8\sqrt{2}"

c.


"x=\dfrac{-4.5+4.5}{2}=0, y=\dfrac{-12.5+13}{2}=0.25"

"M(0,0.25)"

"l=\sqrt{4.5-(-4.5))^2+(13-(-12.5))^2}=7.5\sqrt{13}"


2.

A.

a.


"5x+2y=-10"

"x=0:y=-5"

y-intercept: "(0, -5)"


"y=0:x=-2"

x-intercept: "(-2, 0)"


"y=-2.5(x+2)"

slope "=-2.5"


"y+5=-2.5x"

b.


"13y-2x=3"

"x=0:y=\dfrac{3}{13}"

y-intercept: "(0, \dfrac{3}{13})"


"y=0:x=-\dfrac{3}{2}"

x-intercept: "(-\dfrac{3}{2}, 0)"


"y=\dfrac{2}{13}(x+\dfrac{3}{2})"

slope "=\dfrac{2}{13}"


"y-\dfrac{3}{13}=\dfrac{2}{13}x"

c.


"25y+31x-18=0"

"x=0:y=\dfrac{18}{25}"

y-intercept: "(0, \dfrac{18}{25})"


"y=0:x=\dfrac{18}{31}"

x-intercept: "(\dfrac{18}{31}, 0)"


"y=-\dfrac{31}{25}(x-\dfrac{18}{31})"

slope "=-\dfrac{31}{25}"


"y-\dfrac{18}{25}=-\dfrac{31}{25}x"



d.


"-3x+4y-10=7x-2y+50"

"x=0:y=10"

y-intercept: "(0, 10)"


"y=0:x=-6"

x-intercept: "(-6, 0)"


"y=\dfrac{5}{3}(x-(-6))"

slope "=\dfrac{5}{3}"


"y-10=\dfrac{5}{3}x"

B.

a)


"slope=\dfrac{7-3}{4-0}=1"



b)


"slope=\dfrac{-2-6}{-1-(-9)}=-1"



C.

a)


"y-10=-2x"

"y=-2x+10"

b)


"y-(-2)=-3(x-4)"

"y=-3x+10"

c)


"y=4"

"y=-2x+10"

d)


"slope=\dfrac{1-(-2)}{-12-3}=-\dfrac{1}{5}"


"y-(-2)=-\dfrac{1}{5}(x-3)"

"y=-\dfrac{1}{5}x-\dfrac{7}{5}"

e)


"slope=\dfrac{450-240}{15-20}=-42"


"y-240=-42(x-20)"

"y=-42x+1080"